You have been asked by the president of your company to evaluate the proposed acquisition of a new special-purpose truck for $70,000. The truck falls into the MACRS three-year class, and it will be sold after three years for $5,000. Use of the truck will require an increase in NWC (spare parts inventory) of $10,000. The truck will have no effect on revenues, but it is expected to save the firm $32,000 per year in before-tax operating costs, mainly labor. The firm's marginal tax rate is 40 percent. What will the operating cash flow for this project be during year 2?

MACRS Rates:

Year 1: 33.33%

Year 2: 44.45%

Year 3: 14.81%

Year 4: 7.41%
Answer #1

rate positively ..

Computation of year 2 operating cash flow
i Revenue = 32000
ii Depreciation = 31115
70000*44.45%
iii=i-ii Profit before tax = 885
iv=iii*40% Tax @ 40% 354
v=iii-iv Profit after tax = 531
vi=v+ii Operating cash flow 31646
Ans = 31646